An 18in piece of wire is cut into 3 peices. The third piece is 4 inches more than half the second piece and the first piece is 2 inches shorter than the third piece. Find the length of these three pieces.
Answer by elima(1433)

x = length of the pieces
2nd piece = x
3rd piece = 1/2x+4
1st piece = 3rd piece - 2 = (1/2x+4)-2
Now lets add the pieces together and have them = 18;
x + 1/2x+4 + (1/2x+4)-2 =18
x+ x+4+2=18
2x+6=18
2x=12
x=6
Now we have the length of the 2nd, lets plug this into the others to find their length;
3rd = 1/2(6)+4
3rd = 3+4= 7
1st = 7-2 = 5
check;
5+6=11+7=18
:)